A Carnot refrigerator rejects 2500 kJ of heat at 80°C while using 1100 kJ of work


1
Expert's answer
2021-12-21T12:15:36-0500

Find the heat absorbed:


"Q_a=W+Q_r=1100+2500=3600\text{ kJ}."

The efficiency of the refrigerator is


"\eta=\frac{W}{Q_a}=\frac{1100}{3600}=0.31."